2 ms·
Thinking about this paradox in terms of probabilities: P(probes) = 1 - (1 - P(a civ can launch a probe)) ^ N(civ) If we don't see probes, and if we don't thin
by pklausler 11y ago
Thinking about this paradox in terms of probabilities:
P(probes) = 1 - (1 - P(a civ can launch a probe)) ^ N(civ)
If we don't see probes, and if we don't think P(can launch) is exactly zero, then N(civ) just has to be small.
EDIT: HN deleted my Fortran-style exponentiation.